y = kx(4-x), solve for x
Answer by Jk22(389)

if k<>0 :
y = kx(4-x) => y/k = x(4-x) = 4x - x^2 = -(x-2)^2 + 4

-(x-2)^2 = y/k - 4

(x-2)^2 = 4 - y/k

x-2 = +/-Sqrt(4 - y/k)

x = 2 +/- Sqrt(4 - y/k)


if k=0 : if y=0 any x is solution

if k=0 and y<>0 there are no solution.